# Massless Quasiparticles in Bogoliubov-de Gennes Systems

### v1

@@ G.2 When the $\tilde{\mathcal{P}}$ is anti-unitary
In this case, we only need to consider the case in which $G_{f}(k)$ is unitary (otherwise it reduces to the case with unitary $\tilde{\mathcal{P}}$). A set of reducible minimal nonzero modes are formed by two equivalent representations, namely $D^{\mathbb{R}}(F_{f}(k))\oplus D^{\mathbb{R}}(F_{f}(k))$ both of which belong to the $\mathbb{R}$ class, and a set of irreducible nonzero modes are all irreducible representations of $F_{f}(k)$ which belong to the $\mathbb{H}$ or the $\mathbb{C}$ class. We denote them by $D^{\mathbb{H}}(F_{f}(k))$ and $D^{\mathbb{C}}(F_{f}(k))$ respectively.
For the reducible minimal nonzero modes $D^{\mathbb{R}}(F_{f}(k))\oplus D^{\mathbb{R}}(F_{f}(k))$. Similar to the case where $\tilde{\mathcal{P}}$ is unitary, a ${Z}_{2}$ induction is impossible, but a ${Z}_{2}$ extensions is possible to turn the set of reducible minimal nonzero modes into a set of reducible $\mathcal{C}$-minimal zero modes.
For the irreducible nonzero modes $D^{\mathbb{C}}(F_{f}(k))$. In this case, a $Z_{2}$ induction is possible to turn the set of irreducible nonzero modes into a set of irreducible $\mathcal{C}$-minimal zero modes, which is similar to the case where $\tilde{\mathcal{P}}$ is unitary.
For the irreducible nonzero modes $D^{\mathbb{H}}(F_{f}(k))$ In this case, a single charge operation $\mathcal{C}$ fails to turn the set of irreducible nonzero modes $D^{\mathbb{H}}(F_{f}(k))$ into a set of $\mathcal{C}$-minimal zero modes.
But we can adopt two charge conjugation operators $\mathcal{C}_{1},\mathcal{C}_{2}$ to obtain a set of $\mathcal{C}$-minimal zero modes. In this case we have $F_{f}^{c_{1}\times c_{2}}(k)/F_{f}(k)\cong{Z}_{2}\times{Z}_{2}$, and the representations of the two charge operators $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ should anti-commute with each other
$D(\mathcal{C}_{1})D(\mathcal{C}_{2})=-D(\mathcal{C}_{2})D(\mathcal{C}_{1})$
from the Clifford theoryKarpilovsky [1985].
The process can be explained as an extension followed by an induction. We first do a ${Z}_{2}$ extension by the $\mathcal{C}_{1}$ to make the two irreps contained in $D^{\mathbb{H}}(H_{f}(k))$ to be non-equivalent in $D^{\mathbb{H}}(H^{c_{1}}_{f}(k))$. Hence $\mathcal{C}_{1}$ is represented as $D(\mathcal{C}_{1})=\begin{pmatrix}\mathrm{i}I&0\\ 0&-\mathrm{i}I\end{pmatrix}$. Now the irreducible nonzero mode $D^{\mathbb{H}}(F^{c_{1}}_{f}(k))$ is changed to the $\mathbb{C}$ class. Then we do a ${Z}_{2}$ induction by $\mathcal{C}_{2}$, where $\mathcal{C}_{2}$ is represented as $D(\mathcal{C}_{2})=\begin{pmatrix}0&\mathrm{i}I\\ \mathrm{i}I&0\end{pmatrix}$. After the two steps, a set of irreducible $\mathcal{C}$-minimal zero modes for the group $F_{f}^{c_{1}\times c_{2}}(k)$ are obtained.

### v3

@@ G.2 When the $\tilde{\mathcal{P}}$ is anti-unitary
In this case, we only need to consider the case in which $G_{f}(k)$ is unitary (otherwise it reduces to the case with unitary $\tilde{\mathcal{P}}$). A set of RMNZM are formed by two equivalent representations, namely $\mathcal{D}_{+}^{\mathbb{R}}(F_{f}(k))\oplus\mathcal{D}_{-}^{\mathbb{R}}(F_{f}(k))$ both of which belong to the $\mathbb{R}$ class, and a set of irreducible nonzero modes are all irreps of $F_{f}(k)$ which belong to the $\mathbb{H}$ or the $\mathbb{C}$ class. We denote them by $D^{\mathbb{H}}(F_{f}(k))$ and $D^{\mathbb{C}}(F_{f}(k))$ respectively.
For the RMNZM $\mathcal{D}_{+}^{\mathbb{R}}(F_{f}(k))\oplus\mathcal{D}_{-}^{\mathbb{R}}(F_{f}(k))$. Similar to the case where $\tilde{\mathcal{P}}$ is unitary, a ${Z}_{2}$ induction is impossible, but a ${Z}_{2}$ extensions is possible to turn the set of RMNZM into a set of reducible minimal $\mathcal{C}$-zero modes.
For the irreducible nonzero modes $D^{\mathbb{C}}(F_{f}(k))$. Similar to the case where $\tilde{\mathcal{P}}$ is unitary, a $Z_{2}$ induction is possible to turn the set of irreducible nonzero modes into a set of irreducible $\mathcal{C}$-zero modes.
For the irreducible nonzero modes $D^{\mathbb{H}}(F_{f}(k))$. In this case, a single charge operation $\mathcal{C}$ fails to turn the set of irreducible nonzero modes $D^{\mathbb{H}}(F_{f}(k))$ into $\mathcal{C}$-zero modes.
But we can adopt two charge conjugation operators $\mathcal{C}_{1},\mathcal{C}_{2}$ to obtain a set of $\mathcal{C}$-minimal zero modes. In this case we have $F_{f}^{c_{1}\times c_{2}}(k)/F_{f}(k)\cong{Z}_{2}\times{Z}_{2}$, and the representations of the two charge operators $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ should anti-commute with each other $D(\mathcal{C}_{1})D(\mathcal{C}_{2})=-D(\mathcal{C}_{2})D(\mathcal{C}_{1})$ from the Clifford theoryKarpilovsky [1985].
The process can be understood as an extension followed by an induction. We first do a ${Z}_{2}$ extension by the $\mathcal{C}_{1}$ to make the two irreps contained in $D^{\mathbb{H}}(H_{f}(k))$ to be non-equivalent in $D^{\mathbb{H}}(H^{c_{1}}_{f}(k))$. Hence $\mathcal{C}_{1}$ is represented as $D(\mathcal{C}_{1})=\begin{pmatrix}\mathrm{i}I&0\\ 0&-\mathrm{i}I\end{pmatrix}$. Now the irreducible nonzero mode $D^{\mathbb{H}}(F^{c_{1}}_{f}(k))$ is changed to the $\mathbb{C}$ class. Then we do a ${Z}_{2}$ induction by $\mathcal{C}_{2}$, where $\mathcal{C}_{2}$ is represented as $D(\mathcal{C}_{2})=\begin{pmatrix}0&\mathrm{i}I\\ \mathrm{i}I&0\end{pmatrix}$. After the two steps, a set of irreducible $\mathcal{C}$-zero modes for the group $F_{f}^{c_{1}\times c_{2}}(k)$ are obtained.